# Smoothed Analysis of Online Metric Matching with a Single Sample: Beyond Metric Distortion

### v1

@@ 5 Euclidean Metrics
In this section, we turn to the Euclidean setting, where $\mathcal{X}=[0,1]^{d}$ and $\delta$ is the Euclidean distance. This case is widely studied in online metric matching. We begin by stating our paper’s main theorem, which characterizes the competitive ratio achievable under $\sigma$-smooth request distributions.
@@
We construct a $2^{d}$-ary $2$-HST with height $h$, denoted as $\mathcal{T}$, from $\mathcal{H}$ as follows: each cube in $\mathcal{H}$ corresponds to a node, and the children of a node corresponding to $H\in\mathcal{H}$ are the nodes corresponding to maximal subsets of $H$ in $\mathcal{H}$. For every $x\in[0,1]^{d}$, let $\mathcal{T}(x)$ be the leaf node corresponding to the (unique) cube in $\mathcal{H}_{0}$ that contains $x$. For $S\subseteq[0,1]^{d}$, define $\mathcal{T}(S):=\{\mathcal{T}(s)\mid s\in S\}$. We show in the following lemma that a cost upper bound for an algorithm on the HST metric induced by $\mathcal{T}$ gives rise to a cost upper bound for the corresponding algorithm on $0,1]^{d}$.
@@
Given an algorithm $\mathcal{A}$ for the HST metric induced by $\mathcal{T}$, there exists an algorithm $\mathcal{A}^{\prime}$ for $0,1]^{d}$ such that for all $S$ and $R$, $\mathrm{cost}(\mathcal{A}^{\prime};S,R)\leq\sqrt{d}\cdot(\mathrm{cost}(\mathcal{A};\mathcal{T}(S),\mathcal{T}(R))/4+n\cdot 2^{-h})$.
@@
By Lemma 5.4, it suffices to provide algorithms for a $2^{d}$-ary $2$-HST metric with a certain height $h$, for which we apply Theorems 3.1 and 4.1, where the height $h$ is chosen to minimize the resulting cost for $0,1]^{d}$. We establish upper bounds for the expected cost of the $\mathrm{RS}$ and $\mathrm{BBGN}$ algorithms for the specified HST metrics in the following two corollaries, whose proofs only involve mechanical calculation and are deferred to Appendices A and B, respectively.
@@
For the $2^{d}$-ary $2$-HST metric with height $h:=\lfloor\log(n)/d\rfloor$, and for any $\mathbb{D}=\prod_{i=1}^{n}\mathbb{D}_{i}$,
@@
Lemma 5.7 (Lemma 15 in [YY25]).
Let $\mathbb{D}_{i}$ be a $\sigma$-smooth distribution over $0,1]^{d}$ with $\sigma\in(0,1]$. For all $d\geq 1$ and $S\subseteq[0,1]^{d}$ with $|S|=n$,
$\mathbb{E}_{r_{i}\sim\mathbb{D}_{i}}\left[\min_{s\in S}\left\|{s-{r_{i}}}\right\|_{2}\right]\geq\Omega\left(\sigma^{\frac{1}{d}}n^{-\frac{1}{d}}\right).$
@@
Fix $x\in[0,1-L]$, and we analyze $\mathbb{E}[|\hat{s}(x)-\hat{r}(x)|]$. Note that $\hat{r}(x)\sim\mathrm{PB}(\mathbf{w}(x))$, where $w_{i}(x):=\int_{x}^{x+L}p_{i}(y)\mathrm{d}y$ for each $i\in[n]$. Define $W(x):=\left\|{\mathbf{w}(x)}\right\|_{1}$, which implies $\mathbb{E}[\hat{r}(x)]=W(x)$. The following lemma lower bounds $\mathbb{E}[|\hat{s}(x)-\hat{r}(x)|]$.
@@
$w_{i}(x)=\int_{x}^{x+L}p_{i}(y)\mathrm{d}y\leq\frac{L}{\sigma}\leq\frac{1}{2}.$
@@
$\int_{0}^{1-L}W(x)\mathrm{d}x$ $=\int_{0}^{1-L}\sum_{i=1}^{n}\int_{x}^{x+L}p_{i}(y)\mathrm{d}y\mathrm{d}x=\sum_{i=1}^{n}\int_{0}^{1}p_{i}(y)\int_{0}^{1-L}{\mathbf{1}\left\{{y\in[x,x+L]}\right\}}\mathrm{d}x\mathrm{d}y$
$=\sum_{i=1}^{n}\int_{0}^{1}p_{i}(y)\cdot\min\{L,y,1-y\}\mathrm{d}y\geq L\sum_{i=1}^{n}\int_{L}^{1-L}p_{i}(y)\mathrm{d}y.$
@@
$\int_{L}^{1-L}p_{i}(y)\mathrm{d}y=1-\int_{0}^{L}p_{i}(y)\mathrm{d}y-\int_{1-L}^{1}p_{i}(y)\mathrm{d}y\geq 1-\frac{2L}{\sigma}=\frac{1}{2}.$
As a result,
$\int_{0}^{1-L}W(x)\geq\frac{Ln}{2}=\Omega(\sigma n).$
Since $W(x)\leq n$ for every $x\in[0,1-L]$, we have $\sqrt{W(x)}\geq W(x)/\sqrt{n}$ for every $x\in[0,1-L]$ by the concavity of $f(t)=\sqrt{t}$, which implies that
@@
as desired. ∎
@@ 6 Discussion and Future Directions
In this paper, we study the online metric matching problem for the Euclidean metric $0,1]^{d}$ when servers are adversarial and requests are independently drawn from distinct smooth distributions. We present an $O(1)$-competitive algorithm for $0,1]^{d}$ with $d\neq 2$, given a single sample from each request distribution. A key feature of our approach is that, by directly upper-bounding the algorithm’s cost after a simple deterministic metric embedding, we bypass the $\Omega(\log n)$ competitive-ratio barrier that arises in the adversarial setting due to metric distortion. Since embeddings into HST metrics have already proven extremely effective for related online problems such as $k$-server [BBMN15], $k$-taxi [GKP24], and several variants of online metric matching [EKW16, BFT24], a natural and exciting future direction is to extend our techniques to these problems.
Our guarantees rely on requests being independently sampled. An intriguing direction would be to see what forms of correlation among requests might still permit a $o(\log n)$ competitive ratio. As a starting point, recent breakthroughs in smoothed analysis of online learning [BDGR22, HRS24] allow each arrival’s distribution—while required to be smooth—to depend on the realized history of arrivals and algorithmic decisions, and their techniques may extend to our setting. In addition, the correlation models studied in online stochastic matching [AM23] and prophet inequalities [ISW23] present additional promising avenues.
@@ Acknowledgment
This work was supported in part by NSF grant CCF-2338226.
@@
for the $2^{d}$-ary $2$-HST metric with height $h$. When $d=1$, $H_{2^{d}}=3/2$, and
@@
for the $2^{d}$-ary $2$-HST metric with height $h$. When $d=1$,

### v2

@@ 5 Euclidean Metrics
In this section, we turn to the Euclidean setting, where $\mathcal{X}=[0,1]^{d}$ and $\delta$ is the Euclidean distance. This case is widely studied in online metric matching. We begin by stating our paper’s main theorem, which characterizes the competitive ratio achievable under $\sigma$-smooth request distributions. Recall that a measure $\mu$ over $0,1]^{d}$ is $\sigma$-smooth for $\sigma\in(0,1]$ if $\mu(\mathcal{X})\leq\mathbb{U}(\mathcal{X})/\sigma$ for every measurable subset $\mathcal{X}\subseteq[0,1]^{d}$, where $\mathbb{U}$ is the uniform distribution over $0,1]^{d}$.
@@
We construct a $2^{d}$-ary $2$-HST with height $h$, denoted as $\mathcal{T}$, from $\mathcal{H}$ as follows: each cube in $\mathcal{H}$ corresponds to a node, and the children of a node corresponding to $H\in\mathcal{H}$ are the nodes corresponding to maximal subsets of $H$ in $\mathcal{H}$. For every $x\in[0,1]^{d}$, let $\mathcal{T}(x)$ be the leaf node corresponding to the (unique) cube in $\mathcal{H}_{0}$ that contains $x$. For $S\subseteq[0,1]^{d}$, define $\mathcal{T}(S):=\{\mathcal{T}(s)\mid s\in S\}$. We show in the following lemma that a cost upper bound for an algorithm on $\mathcal{T}$ gives rise to a cost upper bound for the corresponding algorithm on $0,1]^{d}$.
@@
Given an algorithm $\mathcal{A}$ on $\mathcal{T}$, there exists an algorithm $\mathcal{A}^{\prime}$ on $0,1]^{d}$ such that for all $S$ and $R$, $\mathrm{cost}(\mathcal{A}^{\prime};S,R)\leq\sqrt{d}\cdot(\mathrm{cost}(\mathcal{A};\mathcal{T}(S),\mathcal{T}(R))/4+n\cdot 2^{-h})$.
@@
By Lemma 5.4, it suffices to provide algorithms for any $2^{d}$-ary $2$-HST with a certain height $h$, for which we apply Theorems 3.1 and 4.1, where the height $h$ is chosen to minimize the resulting cost for $0,1]^{d}$. We establish upper bounds for the expected cost of the $\mathrm{RS}$ and $\mathrm{BBGN}$ algorithms for the specified HST in the following two corollaries, whose proofs only involve mechanical calculation and are deferred to Appendices A and B, respectively.
@@
For the $2^{d}$-ary $2$-HST with height $h:=\lfloor\log(n)/d\rfloor$, and for any $\mathbb{D}=\prod_{i=1}^{n}\mathbb{D}_{i}$,
@@
Lemma 5.7 (Lemma 14 in the arXiv version of [yang2025online]).
Let $\mathbb{D}_{i}$ be a $\sigma$-smooth distribution over $0,1]^{d}$ with $\sigma\in(0,1]$. For all $d\geq 1$ and finite $S\subseteq[0,1]^{d}$,
$\mathbb{E}_{r_{i}\sim\mathbb{D}_{i}}\left[\min_{s\in S}\left\|{s-{r_{i}}}\right\|_{2}\right]\geq\Omega\left(\sigma^{\frac{1}{d}}|S|^{-\frac{1}{d}}\right).$
@@
Fix $x\in[0,1-L]$, and we analyze $\mathbb{E}[|\hat{s}(x)-\hat{r}(x)|]$. For each $i\in[n]$, we use $\mu_{\mathbb{D}_{i}}$ to denote the density of $\mathbb{D}_{i}$ with respect to the uniform distribution over $[0,1]$. Note that $\hat{r}(x)\sim\mathrm{PB}(\mathbf{w}(x))$, where $w_{i}(x):=\int_{x}^{x+L}\mu_{\mathbb{D}_{i}}(y)\mathrm{d}y$ for each $i\in[n]$. Define $W(x):=\left\|{\mathbf{w}(x)}\right\|_{1}$, which implies $\mathbb{E}[\hat{r}(x)]=W(x)$. The following lemma lower bounds $\mathbb{E}[|\hat{s}(x)-\hat{r}(x)|]$.
@@
$w_{i}(x)=\int_{x}^{x+L}\mu_{\mathbb{D}_{i}}(y)\mathrm{d}y\leq\frac{L}{\sigma}\leq\frac{1}{2}.$
@@
$\int_{0}^{1-L}W(x)\mathrm{d}x$ $=\int_{0}^{1-L}\sum_{i=1}^{n}\int_{x}^{x+L}\mu_{\mathbb{D}_{i}}(y)\mathrm{d}y\mathrm{d}x=\sum_{i=1}^{n}\int_{0}^{1}\mu_{\mathbb{D}_{i}}(y)\int_{0}^{1-L}{\mathbf{1}\left\{{y\in[x,x+L]}\right\}}\mathrm{d}x\mathrm{d}y$
$=\sum_{i=1}^{n}\int_{0}^{1}\mu_{\mathbb{D}_{i}}(y)\cdot\min\{L,y,1-y\}\mathrm{d}y\geq L\sum_{i=1}^{n}\int_{L}^{1-L}\mu_{\mathbb{D}_{i}}(y)\mathrm{d}y.$
@@
$\int_{L}^{1-L}\mu_{\mathbb{D}_{i}}(y)\mathrm{d}y=1-\int_{0}^{L}\mu_{\mathbb{D}_{i}}(y)\mathrm{d}y-\int_{1-L}^{1}\mu_{\mathbb{D}_{i}}(y)\mathrm{d}y\geq 1-\frac{2L}{\sigma}=\frac{1}{2}.$
Combining the above two displayed equations,
$\int_{0}^{1-L}W(x)\geq\frac{Ln}{2}=\Omega(\sigma n).$ (15)
For every $x\in[0,1-L]$, since $W(x)\in[0,n]$, we have $\sqrt{W(x)}\geq W(x)/\sqrt{n}$. It follows that
@@
where the second inequality holds by (15), concluding the proof. ∎
@@ 6 Discussion and Future Directions
In this paper, we study the online metric matching problem for the Euclidean space $0,1]^{d}$ when servers are adversarial and requests are independently drawn from distinct smooth distributions. We present an $O(1)$-competitive algorithm for $0,1]^{d}$ with $d\neq 2$, given a single sample from each request distribution. A key feature of our approach is that, by directly upper-bounding the algorithm’s cost after a simple deterministic metric embedding, we bypass the $\Omega(\log n)$ competitive-ratio barrier that arises in the adversarial setting due to metric distortion. Since metric embeddings into HSTs have already been proven extremely effective for related online problems such as $k$-server [DBLP:journals/jacm/BansalBMN15, DBLP:conf/stoc/BubeckCLLM18], $k$-taxi [DBLP:conf/soda/GuptaKP24], and several variants of online metric matching [DBLP:conf/stoc/EmekKW16, DBLP:conf/soda/BhoreFT24], a natural and exciting future direction is to extend our techniques to these problems.
Our guarantees rely on requests being independently sampled. An intriguing direction would be to see what forms of correlation among requests might still permit an $o(\log n)$ competitive ratio. As a starting point, recent breakthroughs in smoothed analysis of online learning [DBLP:conf/colt/BlockDGR22, DBLP:journals/jacm/HaghtalabRS24] allow each arrival’s distribution—while required to be smooth—to depend on the realized history of arrivals and algorithmic decisions, and their techniques may extend to our setting. In addition, the correlation models studied in online stochastic matching [DBLP:conf/sigecom/AouadM23] and prophet inequalities [DBLP:journals/teco/ImmorlicaSW23] present additional promising avenues.
@@ Acknowledgment
This work was supported in part by NSF grant CCF-2338226. We would like to thank the anonymous reviewers for their many helpful comments and suggestions.
@@
for the $2^{d}$-ary $2$-HST with height $h$. When $d=1$, $H_{2^{d}}=3/2$, and
@@
for the $2^{d}$-ary $2$-HST with height $h$. When $d=1$,